# HaluEval-Wild: Evaluating Hallucinations of Language Models in the Wild

### v2

@@ Abstract
Hallucinations pose a significant challenge to the reliability of large language models (LLMs) in critical domains. Recent benchmarks designed to assess LLM hallucinations within conventional NLP tasks, such as knowledge-intensive question answering (QA) and summarization, are insufficient for capturing the complexities of user-LLM interactions in dynamic, real-world settings. To address this gap, we introduce HaluEval-Wild, the first benchmark specifically designed to evaluate LLM hallucinations in the wild. We meticulously collect challenging (adversarially filtered by Alpaca) user queries from existing real-world user-LLM interaction datasets, including ShareGPT, to evaluate the hallucination rates of various LLMs. Upon analyzing the collected queries, we categorize them into five distinct types, which enables a fine-grained analysis of the types of hallucinations LLMs exhibit, and synthesize the reference answers with the powerful GPT-4 model and retrieval-augmented generation (RAG). Our benchmark offers a novel approach towards enhancing our comprehension and improvement of LLM reliability in scenarios reflective of real-world interactions. Our benchmark is available at https://github.com/Dianezzy/HaluEval-Wild.
HaluEval-Wild:
Evaluating Hallucinations of Language Models in the Wild
Zhiying Zhu Carnegie Mellon University zhiyingz@cs.cmu.edu                        Yiming Yang Carnegie Mellon University yiming@cs.cmu.edu                        Zhiqing Sun Carnegie Mellon University zhiqings@cs.cmu.edu
@@ 1 Introduction
Past hallucination benchmarks have primarily drawn from traditional NLP tasks. Traditionally, researchers have assessed model hallucinations within the confines of machine translation (Zhou et al., 2020), text summarization (Zhao et al., 2020; Qiu et al., 2023), and knowledge-intensive dialogues (Dziri et al., 2022). More recently, attention has shifted towards the evaluation of hallucinations in general-purpose aligned LLMs (Li et al., 2023a, 2024). However, to our knowledge, none have thoroughly evaluated LLM hallucinations in real-world scenarios in the wild.
To bridge this gap, we introduce HaluEval-Wild, the first benchmark designed to assess such general-purpose aligned langauge models “in the wild”. Our approach commenced with an analysis of the ShareGPT dataset, containing over 100,000 dialogues between users and ChatGPT, from which we meticulously filtered to isolate queries that significantly challenge the model’s knowledge and reasoning capabilities. This process involved adversarial filtering against Alpaca (Taori et al., 2023), an elementary-level aligned LLM, to ensure the selection of queries that are difficult enough. This selection process culminated in 500 challenging user queries, categorized into five types. We also use retrieval-augmented generation (Lewis et al., 2020) to produce the reference answers.
@@
2 Related Works
Benchmarking LLM Hallucinations
Internal Knowledge of LLMs
External Knowledge Augmentation
@@ 2 Related Works
Type Query Response
OoS Where am I located? You are located in the United States.
CR Run a CPU stress test from android command line You can run a CPU stress test from the Android command line using the "adb shell" command.
Table 2: Evaluation results across various LLMs. Lower hallucination rates indicate superior performance.
Query Type OoS CR IC BM CE Avg. Avg. Response Rate
GPT-4 Turbo 14.00 33.00 25.25 9.00 12.00 18.64 99.80
GPT-35-Turbo 26.00 60.00 28.28 41.00 22.00 35.47 99.80
Mixtral 8x7B 55.00 60.61 63.27 46.00 33.00 51.51 99.40
Mistral-7B 61.00 69.00 72.45 45.00 40.00 57.43 99.60
Llama 2-Chat 70B 64.00 83.00 34.69 70.71 49.00 60.36 99.40
Llama 2-Chat 13B 48.00 71.72 57.73 61.62 35.00 54.75 99.00
Llama 2-Chat 7B 54.00 73.00 57.73 64.65 33.00 56.45 99.20
Vicuna 13B 48.00 90.00 59.79 60.00 50.00 61.57 99.40
Alpaca 7B 99.00 100.00 100.00 99.00 98.00 99.20 100.00
Table 3: Comparison with popular LLM alignment benchmarks. † reports the result of Alpaca 13B.
Benchmark HaluEval-Wild Avg. $\downarrow$ MT-bench (score) $\uparrow$ AlpacaEval $\uparrow$ AlpacaEval 2.0 $\uparrow$
GPT-4 Turbo 18.64% 9.32 97.70% 50.00%
GPT-35-Turbo 35.47% 8.39 93.42% 14.13%
Mixtral 8x7B 51.51% 8.30 94.78% 18.26%
Mistral-7B 57.43% 6.84 92.78% 14.72%
Llama 2-Chat 70B 60.36% 6.86 92.66% 13.87%
Llama 2-Chat 13B 54.75% 6.65 81.09% 7.70%
Llama 2-Chat 7B 56.45% 6.27 71.37% 4.96%
Vicuna 13B 61.57% 6.39 82.11% 7.14%
Alpaca 7B 99.20%   4.54 26.46% 2.59%
Figure 3: Hallucination rates ($\downarrow$) of direct generation, SR (self-reflection), and hinted SR (hinted self-reflection.
@@ 3 Construction of HaluEval-Wild
Real-user queries are vital for assessing LLM hallucination in practical scenarios. In this context, we introduce HaluEval-Wild, a challenging dataset curated from real-world interactions between individuals and LLMs. The construction pipeline of HaluEval-Wild is shown in Appendix 1.
3.1 Challenging Query Collection
We start with the ShareGPT [^1] https://huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered raw dataset, which contains about 100,000 conversations between users and ChatGPT. Our goal is to identify user queries in ShareGPT that are prone to causing hallucinations. To streamline our approach, we focus specifically on the first round of interactions between users and ChatGPT.
We observe certain common patterns in the well-aligned ChatGPT response (OpenAI, 2022), such as the usage of phrases like "I’m sorry, but" and "As an AI language model," which often indicate that the corresponding query is challenging for the LLM, likely leading to inaccurate responses. Consequently, we label LLM responses containing these patterns as challenging queries prone to inducing hallucinations.
Using the pseudo label, we fine-tune a Llama 2-7B model (Touvron et al., 2023) and configure it to function as an initial classifier tasked with automatically pre-screening challenging queries. This classifier processes the first turn of real-user queries and corresponding LLM responses, generating a binary label that indicates the likelihood of hallucination occurring in the query-induced conversations. The classifier has the potential to capture specific characteristics of user queries beyond mere rule-based or keyword analysis.
@@ 3.2 Fine-grained Categorization
We analyze query-induced hallucinations using a categorization framework, outlining five main types with examples in Appendix 1:
@@
In our investigation, we instruct GPT-4 to automatically categorize 6,505 challenging queries labeled in the previous step (Section 3.1) into the aforementioned five fine-grained categories. In Figure  2, we illustrate the distribution of query types as determined by GPT-4. While acknowledging potential limitations in the precision of GPT-4’s classifications, the presented distribution provides valuable insights into the real-world prevalence of query types prone to inducing hallucinations in LLMs.
To enhance the precision of the GPT-4 classifier, which may be compromised due to its inherent inclination towards hallucination, we undertake manual verification for queries categorized under each type, retaining only those accurately classified and those in which Alpaca exhibits hallucinations. We conduct manual verification until each category reaches 100 instances. This meticulous validation not only confirms the potential for hallucination in such queries but also ensures that these queries pose sufficient challenges for LLMs, like Alpaca, to provide accurate answers. Ultimately, we establish a fine-grained benchmark for hallucination evaluation from user-LLM interactions in the wild.
@@ 3.3 Evaluation with Reference Answers
To facilitate the evaluation of hallucination in LLMs, we provide a reference answer generated by GPT-4 for each user query. To overcome the inherent hallucination challenges of GPT-4 and to provide a proficient response, we incorporate information from an external search engine [^2] https://duckduckgo.com/ by retrieving the top five relevant passages, which are then concatenated with the prompt for GPT-4. With the reference answer, we can evaluate an LLM response by asking GPT-4 to judge whether it is hallucinated. A response is considered non-hallucinatory if it is consistent with the reference answer or if GPT-4 explicitly admits its inability to fulfill the request. The prompts for automatic categorization, reference answer generation and hallucination evaluation are available in Appendix A.
4 Experiments
4.1 Evaluation Models
We evaluate a variety of LLMs on HaluEval-Wild, encompassing both open-source and closed-source models. Open-source models such as Alpaca (7B), Vicuna (13B) Chiang et al. (2023), Llama 2 (7B, 13B, 70B), Mistral (7B), and Mixtral (8x7B) Jiang et al. (2024) are accessed through the vLLM library Kwon et al. (2023). Additionally, we examine closed-source LLMs, including OpenAI’s GPT-4 (Turbo) and GPT-3.5 (Turbo), accessed via Microsoft Azure [^3] https://azure.microsoft.com/en-us/solutions/ai.
@@ 4.2 Main Results & Analysis
We present the evaluation of HaluEval-Wild across various LLMs in Table 2 and Table 3.
@@
As indicated in Table 2, there is a wide variance in hallucination rates among different models when confronted with various types of queries. Alpaca 7B, showing a hallucination rate of 99.20%, underscores a significant challenge in dealing with difficult queries. In contrast, GPT-4 Turbo, with the lowest average hallucination rate of 18.64%, illustrates a superior ability to manage such queries, thereby demonstrating a higher reliability.
HaluEval-Wild vs. Other Benchmarks
@@ 4.3 Hallucination Mitigation with Self-Reflection
We use self-reflection as a representative hallucination mitigation mechanism. Self-reflecion (Shinn et al., 2023; Dhuliawala et al., 2023; Ji et al., 2023) enhances LLM responses effectively by utilizing textual feedback from prior errors. Our experimental setup closely aligns with that of Li et al. (2024) with variations in prompts. We first apply self-reflection with prompts that solely instructed LLMs to correct hallucinations without providing any explicit hints. In the hinted version, we incorporate a description of the hallucination type corresponding to the query type as textual feedback in each iteration.
@@ 5 Conclusion
While HaluEval-Wild offers valuable insights into LLM hallucinations, it is not without its limitations. First, the benchmark’s focus on challenging queries specifically designed to induce hallucinations might not fully encapsulate the breadth of everyday user-LLM interactions. Additionally, the categorization and selection process, despite being rigorous, could introduce biases based on the subjective judgment of what constitutes a challenging query. Furthermore, the reliance on manual verification for categorization accuracy and the generation of reference answers may not capture the full spectrum of potential responses, potentially affecting the benchmark’s generalizability. Lastly, as LLMs continue to evolve rapidly, the static nature of any benchmark, including HaluEval-Wild, means it may not fully represent the capabilities of future models. These limitations underscore the need for continuous updates and refinements to HaluEval-Wild and similar benchmarks, ensuring they remain relevant and effective in assessing LLM performance and reliability.
Appendix A Instructions for GPT-4
@@
Appendix B Benchmark Statistics
Query Type OoS CR IC BM CE
Avg. Query Length 18.94 46.72 32.40 29.45 16.47

### v3

@@ Abstract
Hallucinations pose a significant challenge to the reliability of large language models (LLMs) in critical domains. Recent benchmarks designed to assess LLM hallucinations within conventional NLP tasks, such as knowledge-intensive question answering (QA) and summarization, are insufficient for capturing the complexities of user-LLM interactions in dynamic, real-world settings. To address this gap, we introduce HaluEval-Wild, the first benchmark specifically designed to evaluate LLM hallucinations in the wild. We meticulously collect challenging (adversarially filtered by Alpaca) user queries from ShareGPT, an existing real-world user-LLM interaction datasets, to evaluate the hallucination rates of various LLMs. Upon analyzing the collected queries, we categorize them into five distinct types, which enables a fine-grained analysis of the types of hallucinations LLMs exhibit, and synthesize the reference answers with the powerful GPT-4 model and retrieval-augmented generation (RAG). Our benchmark offers a novel approach towards enhancing our comprehension of and improving LLM reliability in scenarios reflective of real-world interactions. Our benchmark is available at https://github.com/HaluEval-Wild/HaluEval-Wild.
@@ 1 Introduction
Past hallucination benchmarks have primarily drawn from traditional NLP tasks. Traditionally, researchers have assessed model hallucinations within the confines of machine translation (Zhou et al., 2020), text summarization (Zhao et al., 2020; Qiu et al., 2023), and knowledge-intensive dialogues (Dziri et al., 2022). More recently, attention has shifted towards the evaluation of hallucinations in general-purpose aligned LLMs (Li et al., 2023a, 2024). However, to our knowledge, none have thoroughly evaluated LLM hallucinations in real-world scenarios in the wild. (see detailed related works in Appendix A)).
To bridge this gap, we introduce HaluEval-Wild, the first benchmark designed to assess such general-purpose aligned langauge models “in the wild”. As demonstrated in Table  1, HaluEval-Wild is unique in that it captures a broad spectrum of real-world user queries, rather than limiting to specific domains, as compared to existing hallucination benchmarks. Our approach commenced with an analysis of the ShareGPT dataset [^1] https://huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered, containing over 100,000 dialogues between users and ChatGPT, from which we meticulously filtered to isolate queries that significantly challenge the model’s knowledge and reasoning capabilities. This process involved adversarial filtering against Alpaca (Taori et al., 2023), an elementary-level aligned LLM, to ensure the difficulty of selected queries. This selection process culminated in 500 challenging user queries, categorized into five types. We also use retrieval-augmented generation (Lewis et al., 2020) to produce the reference answers.
Dataset Domain(s)
HaluEval-Wild General Domain: Real-World User Queries
HaluEval Wikipedia, KG-based dialogue, Newswire
FACTOOL Wikipedia, Python, Math, Science
HaluEval-2.0 Biomedicine, Finance, Science, Education, Wikipedia
Uhgeval Newswire
Med-halt Medicine
Table 1: Comparison of Hallucination-Detection datasets and their domains.
@@ 2 Construction of HaluEval-Wild
Real-user queries are vital for assessing LLM hallucination in practical scenarios. In this context, we introduce HaluEval-Wild, a challenging dataset curated from real-world interactions between individuals and LLMs. The construction pipeline of HaluEval-Wild is shown in Figure 1.
2.1 Identifying Challenging Queries
Collecting challenging queries serves as an important first step in our pipeline. We start with the ShareGPT raw dataset, which contains about 100,000 multi-turn conversations between users and ChatGPT. Our objective is to identify user queries within ShareGPT that are susceptible to eliciting hallucinations from the language model.
Upon initial examination, we notice certain patterns in well-aligned ChatGPT responses (OpenAI, 2022), such as the use of phrases "I’m sorry, but" and "As an AI language model." These phrases often indicate that the corresponding query is challenging for the LLM, likely resulting in inaccurate responses. Building on our observation, a straightforward approach is to select samples using pattern matching. However, relying solely on fixed patterns would result in a high false negative rate, as it is intractable to exhaustively list all patterns that capture the full spectrum of challenging queries.
To mitigate this limitation, we harness the contextual understanding capabilities of LLMs and develop a classifier based on the Llama-2-7B model. Concretely, we use pattern-matching to coarsely extract 4,270 positive examples. We randomly select an equal number of queries not containing hallucination-prone patterns as negative examples. The combined 8,540 positive and negative examples are used to finetune Llama-2-7B, with a 9:1 train-val ratio. The resulting Llama-based classifer, achieving 81.4% validation accuracy, has desired pattern recognizing ability while overcoming the rigidity of using fixed patterns. (see additional metrics in Appendix E).
We further utilize the classifier to pre-screen all user queries in the ShareGPT dataset, collecting 8,067 queries predicted as positive to form our initial pool of challenging queries.
@@ 2.2 Fine-grained Categorization
Queries in the initial pool are then categorized based on our pre-defined taxonomy for query-induced hallucinations. Namely, we define the categories as follows (see examples in Appendix C):
@@
In our investigation, we instruct GPT-4 to automatically categorize challenging queries labeled in Section 2.1 into the aforementioned five fine-grained categories  [^2] Certain challenging queries do not fit specific categories, so we created an ”Other Types” category. For clarity, we excluded it during evaluation but included it in the final release to provide researchers a more comprehensive resource for studying hallucination phenomena.. In Appendix B, we illustrate the distribution of query types as determined by GPT-4. While acknowledging potential limitations in the precision of GPT-4’s classifications, the presented distribution provides valuable insights into the real-world prevalence of query types prone to inducing hallucinations in LLMs.
To mitigate the impact of GPT-4’s propensity for hallucination on classifier precision, we employ a rigorous validation process. Specifically, we manually verify the classification of queries in each category, only retaining those that are accurately labeled [^3] We employ a 2-round cross-validation approach involving two experts.. Additionally, to ensure the queries are sufficiently challenging, we prompt Alpaca (Taori et al., 2023) with them and filter out those that do not elicit hallucinations from Alpaca. We repeat this process until each category comprises 100 instances. This meticulous validation not only confirms the potential for hallucination in these queries but also guarantees that they pose a sufficient challenge for LLMs to provide accurate responses.
@@ 2.3 Evaluation with Reference Answers
To facilitate the evaluation of hallucination in LLMs, we provide a reference answer generated by GPT-4 for each user query. To overcome the inherent hallucination challenges of GPT-4 and to provide a proficient response, we employ the de facto approach of Retrieval-Augmented Generation (RAG) (Mishra et al., 2024; Wei et al., 2024). Specifically, we incorporate information from an external search engine [^4] https://duckduckgo.com/ by retrieving the top five relevant passages and concatenate them with the prompt provided to GPT-4, thereby improving the accuracy and reliability of the reference answers. With the reference answer, we can evaluate an LLM response by asking GPT-4 to judge whether it is hallucinated. A response is considered non-hallucinatory if it is consistent with the reference answer or if it explicitly admits its inability to fulfill the request. Note, even if the capabilities of SoTA LLMs are constantly evolving, leading to the need of continuously updating the set of reference answers, we believe our pipeline for obtaining reference answers and hallucination evaluation is valid for a longer time horizon. The prompts for automatic categorization, reference answer generation and hallucination evaluation are available in Appendix G. We further validate the effectiveness of GPT-4-as-a-Judge as shown in Appendix F.
Benchmark HaluEval-Wild MT-Bench $\uparrow$ AlpacaEval $\uparrow$ AlpacaEval 2.0 $\uparrow$
OoS $\downarrow$ CR $\downarrow$ IC $\downarrow$ BM $\downarrow$ CE $\downarrow$ Avg. $\downarrow$
GPT-4-Turbo 14.00% 33.00% 25.25% 9.00% 12.00% 18.64% 9.32 97.70% 50.00%
GPT-3.5-Turbo 26.00% 60.00% 28.28% 41.00% 22.00% 35.47% 8.39 93.42% 14.13%
Mixtral 8x7B 55.00% 60.61% 63.27% 46.00% 33.00% 51.51% 8.30 94.78% 18.26%
Mistral 7B 61.00% 69.00% 72.45% 45.00% 40.00% 57.43% 6.84 92.78% 14.72%
Llama-2-Chat 70B 64.00% 83.00% 34.69% 70.71% 49.00% 60.36% 6.86 92.66% 13.87%
Llama-2-Chat 13B 48.00% 71.72% 57.73% 61.62% 35.00% 54.75% 6.65 81.09% 7.70%
Llama-2-Chat 7B 54.00% 73.00% 57.73% 64.65% 33.00% 56.45% 6.27 71.37% 4.96%
Vicuna 13B 48.00% 90.00% 59.79% 60.00% 50.00% 61.57% 6.39 82.11% 7.14%
Alpaca 7B 99.00% 100.00% 100.00% 99.00% 98.00% 99.20%   4.54 26.46% 2.59%
Table 2: Evaluation results across various LLMs on HaluEval-Wild and popular alignment benchmarks. Lower scores on HaluEval-Wild and higher scores on alignment benchmarks indicate superior performance. † reports the result of Alpaca 13B. Note, each model occasionally refuses to respond. Examples without a response are excluded from accuracy calculation. Since the response rate is always above 99%, we omit to report this nuance here.
Figure 2: Hallucination rates ($\downarrow$) of direct generation, SR (self-reflection), and hinted SR (hinted self-reflection).
3 Evaluation and Analysis
@@ 3.1 Main Results & Analysis
We evaluate a variety of LLMs on HaluEval-Wild, encompassing both open-source and closed-source models. As indicated in Table 2, there is a wide variance in hallucination rates among different models when confronted with various types of queries. Alpaca 7B, showing a hallucination rate of 99.20%, underscores a significant challenge in dealing with difficult queries. In contrast, GPT-4-Turbo, with the lowest average hallucination rate of 18.64%, illustrates a superior ability to manage such queries, thereby demonstrating a higher reliability.
HaluEval-Wild vs. Other Alignment Benchmarks
@@
Validating the Effectiveness of RAG in Refining Reference Answers
As shown in Table 2, GPT-4 exhibits a non-trival hallucination rate of 18.64%, which explains why we employed RAG during reference answer generation (discussed in Section 2.3). To further validate the effectiveness of RAG, we randomly select 20 examples from HaluEval-Wild and engage another three experts to assess hallucinations in both the RAG-enhanced GPT-4 responses (i.e., reference answers) and the GPT-4 responses without RAG. Results show that without RAG, GPT-4 has a 20% hallucination rate while it falls to 5% when using RAG. Moreover, experts’ feedbacks indicate that RAG-enhanced GPT-4 responses not only have fewer hallucinations but also excel in providing comprehensive contextual information, thorough consideration of setup parameters, and clear explication of assumptions.
@@ 3.2 Hallucination Mitigation with Self-Reflection
We use self-reflection as a representative hallucination mitigation mechanism. Self-reflecion (Shinn et al., 2023; Dhuliawala et al., 2023; Ji et al., 2023) enhances LLM responses effectively by utilizing textual feedback from prior errors. Our experimental setup closely aligns with that of Li et al. (2024) with variations in prompts. We first apply self-reflection with prompts that only instruct LLMs to correct hallucinations without providing explicit hints. In the hinted version, we incorporate a description of the hallucination type corresponding to the query type as textual feedback in each iteration.
@@ 4 Conclusion
While HaluEval-Wild offers valuable insights into LLM hallucinations, it is not without its limitations. First, the benchmark’s focus on challenging queries specifically designed to induce hallucinations might not fully encapsulate the breadth of everyday user-LLM interactions. Additionally, the categorization and selection process, despite being rigorous, could introduce biases based on the subjective judgment of what constitutes a challenging query. Furthermore, the reliance on manual verification for categorization accuracy and the generation of reference answers may not capture the full spectrum of potential responses, potentially affecting the benchmark’s generalizability. Moreover, although RAG is still the common practice to improve the factuality and reduce hallucinations in LLMs, it potentially induces the recall and faithfulness problems (Zhou et al., 2023; Xie et al., 2023; Yu et al., 2023). Lastly, as LLMs continue to evolve rapidly, the static nature of any benchmark, including HaluEval-Wild, means it may not fully represent the capabilities of future models. These limitations underscore the need for continuous updates and refinements to HaluEval-Wild and similar benchmarks, ensuring they remain relevant and effective in assessing LLM performance and reliability.
Ethics Statement
We are committed to maintaining strong ethical standards in the creation and use of our benchmark dataset. To ensure the privacy and security of individuals, we use the manual verification step to confirm that the dataset does not contain any personally identifiable information (PII). We meticulously reviewed all data entries, removing any information that could potentially identify individuals. By taking these measures, we aim to protect the privacy of individuals and adhere to ethical guidelines in data collection and dissemination.
Appendix A Related Works
Benchmarking LLM Hallucinations
Internal Knowledge of LLMs
External Knowledge Augmentation
Appendix B Distribution of Query Types
Appendix C Examples of Various Categories of Challenging Queries.
Type Query Response
OoS Where am I located? You are located in the United States.
[…]
Appendix D Benchmark Statistics
Query Type OoS CR IC BM CE
Avg. Query Length 18.94 46.72 32.40 29.45 16.47
Appendix E Additional Metrics of the Challenging Query Classifier
Precision Recall F1-score
Positive 83% 78% 81%
Negative 80% 84% 82%
Table 5: Precison, Recall, and F1-score of the Llama-2-based Challenging Query Classifier.
Appendix F GPT-4-as-a-Judge vs. Human-as-a-Judge
[…]
The noted discrepancy may be attributed to the fact that GPT-3.5-turbo and GPT-4-turbo are the two most powerful models listed in Table  2. More advanced models tend to produce responses that are contextually sophisticated, which poses a challenge in identifying hallucinations.
Model TN FP FN TP
GPT-3.5-Turbo 14 1 2 3
GPT-4-Turbo 2 13 4 2 1
Table 6: Confusion matrices for GPT-4-as-a-Judge and Human-as-a-Judge in evaluating GPT-3.5-turbo and GPT-4-turbo responses. Here, ’positive’ indicates the presence of hallucinations in the LLM response.
Appendix G Instructions for GPT-4